40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 3 22:41:17 2019, shruti, Update, General, PLL for PM measurement  PM_measurement.jpeg
    Reply  Wed Dec 4 20:32:43 2019, gautam, Update, General, PLL for PM measurement  
    Reply  Mon Dec 9 21:22:46 2019, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT 
       Reply  Tue Dec 17 20:08:09 2019, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT 20191217.png
          Reply  Thu Jan 16 19:32:23 2020, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT 
             Reply  Thu Jan 23 20:08:49 2020, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT Image-1.jpg
                Reply  Tue Jan 28 19:40:15 2020, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT PMTF.pdfBeatSpectrum.pdf
Message ID: 15088     Entry time: Mon Dec 9 21:22:46 2019     In reply to: 15069     Reply to this: 15101
Author: shruti 
Type: Update 
Category: General 
Subject: PLL / PM measurement of Xend NPRO PZT 

In short:

Using the same setup as before with a LPF changed to have a cutoff of 5 MHz, the PLL was implemented and a TF measurement of the phase modulation was attempted. But, the beatnote drift was too high to get a prolonged phase lock (many times over 5MHz in <5 min).
 

Steps undertaken:

1. Normally I would unlock the IMC (Disabling the servo between the 'Filter' and 'Polarity' on the Mode Cleaner Servo Screen), but today I did not have to since Rana had kept it unlocked.

2. Misaligned the ITMX. This is to prevent cavity resonances from returning to the laser

3. Turned up the air on the HEPA at the PSL table to 100% during the measurement

4. Cables were connected as before (diagram shown in attachment of elog 15069)

5. The X end laser NPRO was actuated for the TF measurement using a long cable connected to TO AUX_X LASER PZT

 

Thoughts and observations:

- Reading out the error signal after amplification cannot distinguish between a locked loop or one out of its range. The error signal would be very small in both cases.

- Looking at the beat note on an oscilloscope, there also seemed to be an additional amplitude modulation that I had not noticed earlier. Rana suggested that it may have something to do with the pre-mode cleaner and the AOM being driven at 80 MHz

- Even though the TF was attempted, it seemed too noisy, suggesting that the PLL did not seem to work

- Rana also suggested that it may be a better idea to use the PZT of one of the lasers as the VCO for the PLL feedback instead of the Marconi.

 

Quote:

I worked on the setup up for the phase modulation measurement of the X end NPRO PZT. A previous similar measurement can be found here (12077). The setup was assembled based on the schematic in Attachment1.

Mixer used: Level 7, Mini circuits ZP-3+
LPF: up to 1.9MHz


Cables exiting the PSL table:
1. LO (Marconi -> Mixer)
2. RF (PSL+X beat note -> Mixer) The cable for this was taken from the Beat Mouth (otherwise connected to the oscilloscope)
3. Ext modulator (SR560 -> Marconi)

The long cable labled 'X Green Beat' was used to connect to the PZT (from the network analyzer).

Observations: The beat note kept floating between 0 and ~100 MHz

The PLL part of the circuit was tested coarsely with the spectrum analyzer function of the Agilent, where the loop was seen to stabilize when the carrier frequency of the Marconi was close to the instantaneous beat frequency.

 

 

 

ELOG V3.1.3-